# arXiv:2603.00183v1[cs.SE] 26 Feb 2026

### v1

@@ 2.3 Foundational works on TCP
To the best of our knowledge, the first mention of TCP occurs in the work of Wong et al. [1997], where it is used as an additional step after test case selection. The first comprehensive look at TCP was then provided by Rothermel et al. [1999], where a wide range of prioritization methods were investigated.
The test case prioritization problem was previously formally defined in the work of Rothermel et al. [2001], and the definition can be seen in Equation˜1:
@@
Over the years, multiple TCP approach categorizations have been proposed. Examples include the taxonomies used by Khatibsyarbini et al. [2018], Singh et al. [2012], and Mukherjee and Patnaik [2021]. For the purpose of the study, the classification introduced by Yoo and Harman [2012] was selected. It includes the following categories of TCP approaches Yoo and Harman [2012]: coverage-based, distribution-based, human-based, probabilistic, history-based, requirements-based, model-based, cost-aware, and other approaches. Using insights collected during the SR, it can be observed that machine learning solutions are underrepresented in the list of approaches. Pan et al. [2021] provided a separate taxonomy for just ML-based approaches, with the following categories: supervised learning, unsupervised learning, reinforcement learning, and natural language processing. Both described classifications are presented side by side in Figure˜4.
Figure 4: Taxonomies of TCP (on a basis of Yoo and Harman [2012]) and ML-based TCP (on a basis of Pan et al. [2021]) approaches.
In addition to the taxonomies in Figure˜4, the TCP approaches are characterized by their properties. Most existing TCP methods can be classified as either static or dynamic Luo et al. [2016]. Static techniques operate only on the source and test code, using only static program analysis. However, dynamic techniques utilize rich run-time semantics. This enables the algorithm to dynamically reorder the cases as the tests are executed. The techniques can also be divided into black-box and white-box types Henard et al. [2016]. Black-box approaches do not require access to the source code of the system under test; in contrast, white-box approaches can use the source code information to improve the prioritization performance.
@@ 2.4 Approach combinators
After an analysis of the titles and full abstracts of all publications identified during the undertaken snowballing literature review, only two works adjacent to the methods proposed in the thesis were identified.
Huang et al. [2021] attempted to combine the results of multiple heuristic TCP methods using ML algorithms. They focused on model-based prioritization and collected six simple estimators rating test cases based on state transitions. The outputs from particular heuristics are then accumulated using a random forest model to produce the final prediction. Their research differs from the approach we propose in this paper in the following ways: it focuses only on model-based TCP, it works only on top of certain methods, and it requires prior training.
Hassan et al. [2024] investigated the use of random sorting to resolve the ties in the ranking of the test cases. They concluded that this method does not yield optimal results and should only be used as a final resort. The approach proposed in this paper supports random tie-breaking, but it also offers more sophisticated strategies.
An additional ad hoc search was performed, after the solutions from this thesis had already been implemented, in an attempt to find similar works in the literature. Two additional articles, described below, were identified.
The research of Cao et al. [2022] used a simple voting scheme to combine the results of multiple models. Each model assigned votes in a decreasing sequence, and the final score was calculated as the sum of the votes and used to determine the final result. Their research differs from the proposed model in the following ways: they focused only on user interface testing, the combination used the simplest voting scheme available, and no method to assign different weights to models was presented.
The work of Mondal and Nasre [2021], who introduced the Hansie framework for TCP, is most adjacent to our proposed solutions. They assumed that internal prioritizer mechanisms are unknown. Then, the Borda count is used for result combination, similarly to one of the proposed approach combinators. Despite resemblances, we believe that their technique is not equivalent. The Borda count algorithm constitutes only a minor part of the research conducted, and Hansie uses parallel prioritization to resolve ties, while the proposal introduces different methods to deal with ties.
@@ 2.5 Metrics for TCP evaluation
Rothermel et al. [1999] introduced the average of the percentage of faults detected (APFD) metric in 1999 Lou et al. [2019], Zhao et al. [2023]. Although it exhibits multiple limitations, it is still the most widely used TCP metric in the literature Abubakar et al. [2023], de S. Campos Junior et al. [2017], Catal and Mishra [2013], Singh et al. [2012], Khatibsyarbini et al. [2018], Mukherjee and Patnaik [2021]. APFD measures how quickly an ordered test case list detects faults. The metric ranges from 0 to 1, with higher values of the metric representing better orderings Elbaum et al. [2000], Catal and Mishra [2013], Mukherjee and Patnaik [2021], Zhao et al. [2023], Lou et al. [2019]. It can be calculated using the formula from Equation˜2 Lou et al. [2019], Wang et al. [2020], Zhao et al. [2023]:
@@
The cost-cognizant weighted average percentage of faults detected ($\text{APFD}_{C}$) introduced by Elbaum et al. [2001] extends APFD to incorporate differences in fault severities and test case costs (for example, their execution times). The $\text{APFD}_{C}$ metric is very popular and is sometimes considered the second most widely used TCP metric de S. Campos Junior et al. [2017], Singh et al. [2012]. It can be seen as an extension of APFD, where test case costs and fault severities are known, and can be calculated as shown in Equation˜3 Wang et al. [2020], Lou et al. [2019]:
@@
where $s$, $n$, $m$, $\text{TF}_{j}$ have the same meaning as in Equation˜2 and $f_{i}$ and $t_{j}$ describe the severity of $i$-th fault and execution cost of $j$-th test case, respectively. It has the same value range as APFD Wang et al. [2020]. In real-world studies, fault severities are rarely known. The $\text{APFD}_{C}$ formula also works when all severities are assumed to be equal Elbaum et al. [2001], Wang et al. [2020], Lou et al. [2019].
Whenever we introduce constraints to TCP, some faults might not get detected, as some test cases are not run. Qu et al. [2007] introduced normalized APFD (NAPFD) in 2007, including an additional $p$ factor that represents the fraction of identified faults in all identifiable faults Abubakar et al. [2023], Lou et al. [2019], Wang et al. [2020], Zhao et al. [2023]. The formula for NAPFD is given in Equation˜4, where $\text{TF}_{j}$ is set to zero if a fault is never detected Lou et al. [2019], Wang et al. [2020]:
@@
Zhao et al. [2023] identified weaknesses in existing metrics from the APFD family, namely that APFD has different value ranges depending on fault and test case counts. This also means that although it returns results bounded by 0 and 1, the optimal solution does not have an APFD of 1, and neither does the worst solution have an APFD of 0. They addressed this weakness, min-max mapping the measure, introducing the rectified APFD ($r\text{APFD}$), which ensures that the optimal solution has $r\text{APFD}=1$. The value of $r\text{APFD}$ for a given test suite is shown in Equation˜5:
@@
Bertolino et al. [2020] introduced the rank percentile average (RPA) and its normalization called normalized-rank-percentile-average (NRPA) to evaluate the order of the test suite independently of the algorithm and the choice of criteria Zhao et al. [2023], Bagherzadeh et al. [2022]. These measures have been criticized before Zhao et al. [2023], Bagherzadeh et al. [2022]. Bagherzadeh et al. [2022] note that NRPA treats successful and failed test cases equally, while failures are generally more important. Zhao et al. [2023] support the limitation described previously and additionally report that sometimes worse sequences report higher NRPA values.
The applicability of TCP is also evaluated in the literature. If a method has a high cost, it might be really effective, but not applicable in practical contexts. Various time-related metrics are most commonly used for this goal Khatibsyarbini et al. [2018], Prado Lima and Vergilio [2020]. Prado Lima and Vergilio [2022] defined the normalized time reduction (NTR) metric to measure time savings gained if the cycle stopped at the first detected fault. It can be determined using Equation˜6 Prado Lima and Vergilio [2022], Zhao et al. [2023]:
@@
where $\text{CI}^{\text{fail}}$ is the number of failed cycles, $\hat{r}_{i}$ is the execution time of $i$-th failing cycle, and $r_{i}$ is the time to first fault of $i$-th failing cycle. The NTR measure was later adopted by Zhao et al. [2023] to evaluate the applicability of TCP models.
All previously described measures are compared in Table˜3.
@@
APFD Rothermel et al. [1999] failed cycles
$\text{APFD}_{C}$ Elbaum et al. [2001] failed cycles
@@
NAPFD Qu et al. [2007] failed cycles
$r\text{APFD}$ Zhao et al. [2023] failed cycles
RPA Bertolino et al. [2020] all cycles
NRPA Bertolino et al. [2020] all cycles
NTR Prado Lima and Vergilio [2022] entire history
@@ 3.1 Research infrastructure (TCPFramework)
Although many works in the field of TCP publish associated software packages to aid reproducible research Mondal and Nasre [2021], Zhao et al. [2023], Bertolino et al. [2020], Luo et al. [2016], Henard et al. [2016], most of them are ad hoc scripts used to collect results for research questions. We were unable to find a comprehensive and unified framework for the development and evaluation of TCP approaches in the existing literature. To aid further research, both for the scope of this research and for other researchers, we propose an extensible platform for the development of test case prioritization methods, called TCPFramework.
@@
  • 1.
    BaseOrder: The base order approach executes the test cases from the suite in their original arrangement Rothermel et al. [1999]. This method can be used as a "no prioritization performed" baseline to evaluate the gains of applying TCP to a certain software project.
  • 2.
    RandomOrder: The random approach unmethodically shuffles the test suite and executes test cases in the resulting order Rothermel et al. [1999].
Other methods are history-based and rate test cases using aggregations of their previous performances. Exponential smoothing was also previously used to aggregate historical results for use in TCP Kim and Porter [2002], Mattis et al. [2020]. It is defined using the formula from Equation˜7:
@@
  • 1.
    FoldFailsOrder: The failure folding approach applies an aggregation function to failure counts of particular test cases from each subsequent cycle. The most notable folders include summation and exponential smoothing. Demonstrated fault effectiveness (DFE), previously used by Kim and Porter [2002] is a special case of this approach.
  • 2.
@@ 3.2 Code representation approach (CodeDistOrder)
The next developed approach, CodeDistOrder, builds on the work of Yang et al. [2023] and Zhou et al. [2024]. Code representation aims to embed the rich information present in the source of a program as a numeric vector Yang et al. [2023], Hrkút et al. [2023]. While the work of Yang et al. [2023] addressed information retrieval-based test prioritization and Zhou et al. [2024] used code representation to cluster test cases, the approach presented in this paper instead focuses on the distances between the code representation vectors. These distance measures include the Manhattan distance, the Euclidean distance, and cosine similarity Hrkút et al. [2023].
The goal of the proposed optimizer is to find the longest path, in terms of the source code representation distance, going through all test cases. Finding an exact solution to such a problem is computationally difficult Ledru et al. [2012]. Instead, we employ a greedy algorithm that selects the highest possible distance at each step. The proposed method goes through the following steps:
  1. 1.
    The code representation is calculated for each case using a vectorizer.
  2. 2.
    Representation distances are calculated between all unique pairs of cases.
  3. 3.
    The starting point is selected as the test case that has the highest aggregated distance from all others in the suite. The prioritized set is initialized to contain only the starting point, and the remaining set is set to all test cases with the exception of the starting point.
  4. 4.
    While the remaining set is nonempty: The test case with the highest aggregated distance from the prioritized set is selected and executed. The test case is then removed from the remaining set and inserted into the prioritized set.
The approach can be tuned using numerous parameters:
  • 1.
    Normalization: The vectors produced contain information on both the semantics and syntax of the source code Yang et al. [2023]. We argue that the syntactic style does not affect the fault-detecting capabilities of test cases. To minimize its influence, we tokenize the source code and remove all unnecessary whitespace.
  • 2.
    Distance: The cosine similarity, Euclidean distance, and Manhattan distance are available as distance measures.
  • 3.
    Aggregation: The distance between a vector and a set of vectors is ambiguous. We support minimal, average, and maximal distances, as used in the literature Ledru et al. [2012].
  • 4.
    Failure adaptation: We predict that once a fault is located, it is more beneficial to focus on test cases similar to it rather than continuing the diversity-focused search. The failure adaptation mechanism switches the optimizer from maximizing the distance to minimizing for the next few searches if the previous test case failed.
@@ 3.4 Mixers
Mixers take multiple prioritization algorithms and mix their weighted results, returning the final test case sequence. The primary intuition is that various test case properties should be incorporated into an algorithm to produce high-quality predictions. New approaches are produced, which for each cycle prioritize the test suite independently using each sub-approach and then merge the results, called queues, into one final list using some algorithm. No internal information leaks from any sub-algorithm. The process is shown in Figure˜5.
@@
The only remaining detail is the way the results are merged. The simplest proposed approach, further called RandomMixedOrder takes the first test case from the $i$-th queue, where $i$ was chosen randomly according to the specified weights. Other proposed merging schemes come directly from electoral system theory. Firstly, BordaMixedOrder uses the Borda method for positional voting to arrange the suite Brandt et al. [2016]. For each sub-approach outcome, we assign points to test cases based on their position. The first case in a given queue gets n1n-1 points (where $n$ is the size of the suite), the second gets n2n-2, etc. until the penultimate and last test cases get 1 and 0 points, respectively. If two test cases are tied, their points are averaged. The points are then multiplied by the weight of a given sub-method and summed up over all queues for each case. The final order is determined by the sum of the scores of all the cases. The Schulze method is also implemented, as SchulzeMixedOrder. It is much more robust than the Borda voting system, providing certain desirable qualities, such as the majority winner rule – if a test case has more than 50% points, it must win Brandt et al. [2016]. It should be noted that, as a variant of the Floyd-Warshall algorithm is used, the computational complexity of this combinator reaches $O\left(n^{3}\right)$, making it unsuitable for larger collections of test cases.

### v2

@@ 2.3 Foundational works on TCP
To the best of our knowledge, the first mention of TCP occurs in the work of Wong et al. (1997), where it is used as an additional step after test case selection. The first comprehensive look at TCP was then provided by Rothermel et al. (1999), where a wide range of prioritization methods were investigated.
The test case prioritization problem was previously formally defined in the work of Rothermel et al. (2001), and the definition can be seen in Equation˜1:
@@
Over the years, multiple TCP approach categorizations have been proposed. Examples include the taxonomies used by Khatibsyarbini et al. (2018), Singh et al. (2012), and Mukherjee and Patnaik (2021). For the purpose of the study, the classification introduced by Yoo and Harman (2012) was selected. It includes the following categories of TCP approaches (Yoo and Harman, 2012): coverage-based, distribution-based, human-based, probabilistic, history-based, requirements-based, model-based, cost-aware, and other approaches. Using insights collected during the SR, it can be observed that machine learning solutions are underrepresented in the list of approaches. Pan et al. (2021) provided a separate taxonomy for just ML-based approaches, with the following categories: supervised learning, unsupervised learning, reinforcement learning, and natural language processing. Both described classifications are presented side by side in Figure˜4.
Figure 4: Taxonomies of TCP (on a basis of (Yoo and Harman, 2012)) and ML-based TCP (on a basis of (Pan et al., 2021)) approaches.
In addition to the taxonomies in Figure˜4, the TCP approaches are characterized by their properties. Most existing TCP methods can be classified as either static or dynamic (Luo et al., 2016). Static techniques operate only on the source and test code, using only static program analysis. However, dynamic techniques utilize rich run-time semantics. This enables the algorithm to dynamically reorder the cases as the tests are executed. The techniques can also be divided into black-box and white-box types (Henard et al., 2016). Black-box approaches do not require access to the source code of the system under test; in contrast, white-box approaches can use the source code information to improve the prioritization performance.
@@ 2.4 Approach combinators
After an analysis of the titles and full abstracts of all publications identified during the undertaken systematic review, only two works adjacent to the methods proposed in the work were identified.
Huang et al. (2021) attempted to combine the results of multiple heuristic TCP methods using ML algorithms. They focused on model-based prioritization and collected six simple estimators rating test cases based on state transitions. The outputs from particular heuristics are then accumulated using a random forest model to produce the final prediction. Their research differs from the approach we propose in this paper in the following ways: it focuses only on model-based TCP, it works only on top of certain methods, and it requires prior training.
Hassan et al. (2024) investigated the use of random sorting to resolve the ties in the ranking of the test cases. They concluded that this method does not yield optimal results and should only be used as a final resort. The approach proposed in this paper supports random tie-breaking, but it also offers more sophisticated strategies.
An additional ad hoc search was performed, after the solutions from this work had already been implemented, in an attempt to find similar works in the literature. Two additional articles, described below, were identified.
The research of Cao et al. (2022) used a simple voting scheme to combine the results of multiple models. Each model assigned votes in a decreasing sequence, and the final score was calculated as the sum of the votes and used to determine the final result. Their research differs from the proposed model in the following ways: they focused only on user interface testing, the combination used the simplest voting scheme available, and no method to assign different weights to models was presented.
The work of Mondal and Nasre (2021), who introduced the Hansie framework for TCP, is most adjacent to our proposed solutions. They assumed that internal prioritizer mechanisms are unknown. Then, the Borda count is used for result combination, similarly to one of the proposed approach combinators. Despite resemblances, we believe that their technique is not equivalent. The Borda count algorithm constitutes only a minor part of the research conducted, and Hansie uses parallel prioritization to resolve ties, while the proposal introduces different methods to deal with ties.
@@ 2.5 Metrics for TCP evaluation
Rothermel et al. (1999) introduced the average of the percentage of faults detected (APFD) metric in 1999 (Lou et al., 2019; Zhao et al., 2023). Although it exhibits multiple limitations, it is still the most widely used TCP metric in the literature (Abubakar et al., 2023; de S. Campos Junior et al., 2017; Catal and Mishra, 2013; Singh et al., 2012; Khatibsyarbini et al., 2018; Mukherjee and Patnaik, 2021). APFD measures how quickly an ordered test case list detects faults. The metric ranges from 0 to 1, with higher values of the metric representing better orderings (Elbaum et al., 2000; Catal and Mishra, 2013; Mukherjee and Patnaik, 2021; Zhao et al., 2023; Lou et al., 2019). It can be calculated using the formula from Equation˜2 (Lou et al., 2019; Wang et al., 2020; Zhao et al., 2023):
@@
The cost-cognizant weighted average percentage of faults detected ($\text{APFD}_{C}$) introduced by Elbaum et al. (2001) extends APFD to incorporate differences in fault severities and test case costs (for example, their execution times). The $\text{APFD}_{C}$ metric is very popular and is sometimes considered the second most widely used TCP metric (de S. Campos Junior et al., 2017; Singh et al., 2012). It can be seen as an extension of APFD, where test case costs and fault severities are known, and can be calculated as shown in Equation˜3 (Wang et al., 2020; Lou et al., 2019):
@@
where $s$, $n$, $m$, $\text{TF}_{j}$ have the same meaning as in Equation˜2 and $f_{i}$ and $t_{j}$ describe the severity of $i$-th fault and execution cost of $j$-th test case, respectively. It has the same value range as APFD (Wang et al., 2020). In real-world studies, fault severities are rarely known. The $\text{APFD}_{C}$ formula also works when all severities are assumed to be equal (Elbaum et al., 2001; Wang et al., 2020; Lou et al., 2019).
Whenever we introduce constraints to TCP, some faults might not get detected, as some test cases are not run. Qu et al. (2007) introduced normalized APFD (NAPFD) in 2007, including an additional $p$ factor that represents the fraction of identified faults in all identifiable faults (Abubakar et al., 2023; Lou et al., 2019; Wang et al., 2020; Zhao et al., 2023). The formula for NAPFD is given in Equation˜4, where $\text{TF}_{j}$ is set to zero if a fault is never detected (Lou et al., 2019; Wang et al., 2020):
@@
Zhao et al. (2023) identified weaknesses in existing metrics from the APFD family, namely that APFD has different value ranges depending on fault and test case counts. This also means that although it returns results bounded by 0 and 1, the optimal solution does not have an APFD of 1, and neither does the worst solution have an APFD of 0. They addressed this weakness, min-max mapping the measure, introducing the rectified APFD ($r\text{APFD}$), which ensures that the optimal solution has $r\text{APFD}=1$. The value of $r\text{APFD}$ for a given test suite is shown in Equation˜5:
@@
Bertolino et al. (2020) introduced the rank percentile average (RPA) and its normalization called normalized-rank-percentile-average (NRPA) to evaluate the order of the test suite independently of the algorithm and the choice of criteria (Zhao et al., 2023; Bagherzadeh et al., 2022). These measures have been criticized before (Zhao et al., 2023; Bagherzadeh et al., 2022). Bagherzadeh et al. (2022) note that NRPA treats successful and failed test cases equally, while failures are generally more important. Zhao et al. (2023) support the limitation described previously and additionally report that sometimes worse sequences report higher NRPA values.
The applicability of TCP is also evaluated in the literature. If a method has a high cost, it might be really effective, but not applicable in practical contexts. Various time-related metrics are most commonly used for this goal (Khatibsyarbini et al., 2018; Prado Lima and Vergilio, 2020). Prado Lima and Vergilio (2022) defined the normalized time reduction (NTR) metric to measure time savings gained if the cycle stopped at the first detected fault. It can be determined using Equation˜6 (Prado Lima and Vergilio, 2022; Zhao et al., 2023):
@@
where $\text{CI}^{\text{fail}}$ is the number of failed cycles, $\hat{r}_{i}$ is the execution time of $i$-th failing cycle, and $r_{i}$ is the time to first fault of $i$-th failing cycle. The NTR measure was later adopted by Zhao et al. (2023) to evaluate the applicability of TCP models.
All previously described measures are compared in Table˜4.
@@
APFD (Rothermel et al., 1999) failed cycles
$\text{APFD}_{C}$ (Elbaum et al., 2001) failed cycles
@@
NAPFD (Qu et al., 2007) failed cycles
$r\text{APFD}$ (Zhao et al., 2023) failed cycles
RPA (Bertolino et al., 2020) all cycles
NRPA (Bertolino et al., 2020) all cycles
NTR (Prado Lima and Vergilio, 2022) entire history
@@ 3.1 Research infrastructure (TCPFramework)
Although many works in the field of TCP publish associated software packages to aid reproducible research (Mondal and Nasre, 2021; Zhao et al., 2023; Bertolino et al., 2020; Luo et al., 2016; Henard et al., 2016), most of them are ad hoc scripts used to collect results for research questions. We were unable to find a comprehensive and unified framework for the development and evaluation of TCP approaches in the existing literature. To aid further research, both for the scope of this research and for other researchers, we propose an extensible platform for the development of test case prioritization methods, called TCPFramework.
@@
  • BaseOrder: The base order approach executes the test cases from the suite in their original arrangement (Rothermel et al., 1999). This method can be used as a "no prioritization performed" baseline to evaluate the gains of applying TCP to a certain software project.
  • RandomOrder: The random approach unmethodically shuffles the test suite and executes test cases in the resulting order (Rothermel et al., 1999).
  • RecentnessOrder: The recentness order rates test cases on the basis of how recently they were introduced to the project. A counter is kept for each case, accumulating the times it was seen in a cycle. The test cases that appeared fewer times in previous regression testing cycles are executed first. This reflects the intuition that verdicts of new test cases are less predictable (since less historical information can be used to estimate their outputs).
Other methods are history-based and rate test cases using aggregations of their previous performances. Exponential smoothing was also previously used to aggregate historical results for use in TCP (Kim and Porter, 2002; Mattis et al., 2020). It is defined using the formula from Equation˜7:
@@
  • FoldFailsOrder: The failure folding approach applies an aggregation function to failure counts of particular test cases from each subsequent cycle. The most notable folders include summation and exponential smoothing. Demonstrated fault effectiveness (DFE), previously used by Kim and Porter (2002) is a special case of this approach.
@@ 3.2 Code representation approach (CodeDistOrder)
The next developed approach, CodeDistOrder, builds on the work of Yang et al. (2023) and Zhou et al. (2024). Code representation aims to embed the rich information present in the source of a program as a numeric vector (Yang et al., 2023; Hrkút et al., 2023). While the work of Yang et al. (2023) addressed information retrieval-based test prioritization and Zhou et al. (2024) used code representation to cluster test cases, the approach presented in this paper instead focuses on the distances between the code representation vectors. These distance measures include the Manhattan distance, the Euclidean distance, and cosine similarity (Hrkút et al., 2023).
The goal of the proposed optimizer is to find the longest path, in terms of the source code representation distance, going through all test cases. Finding an exact solution to such a problem is computationally difficult (Ledru et al., 2012). Instead, we employ a greedy algorithm that selects the highest possible distance at each step.
@@ 3.4 Mixers
Mixers take multiple prioritization algorithms and mix their weighted results, returning the final test case sequence. The primary intuition is that various test case properties should be incorporated into an algorithm to produce high-quality predictions. New approaches are produced, which for each cycle prioritize the test suite independently using each sub-approach and then merge the results, called queues, into one final list using some algorithm. No internal information leaks from any sub-algorithm. We illustrate our process in Figure˜5.
@@
The only remaining detail is the way the results are merged. The simplest proposed approach, further called RandomMixedOrder takes the first test case from the $i$-th queue, where $i$ was chosen randomly according to the specified weights. Other proposed merging schemes come directly from electoral system theory. Firstly, BordaMixedOrder uses the Borda method for positional voting to arrange the suite (Brandt et al., 2016). For each sub-approach outcome, we assign points to test cases based on their position. The first case in a given queue gets n1n-1 points (where $n$ is the size of the suite), the second gets n2n-2, etc. until the penultimate and last test cases get 1 and 0 points, respectively. If two test cases are tied, their points are averaged. The points are then multiplied by the weight of a given sub-method and summed up over all queues for each case. The final order is determined by the sum of the scores of all the cases. The Schulze method is also implemented, as SchulzeMixedOrder. It is much more robust than the Borda voting system, providing certain desirable qualities, such as the majority winner rule – if a test case has more than 50% points, it must win (Brandt et al., 2016). It should be noted that, as a variant of the Floyd-Warshall algorithm is used, the computational complexity of this combinator reaches $O\left(n^{3}\right)$, making it unsuitable for larger collections of test cases.
We selected these three aggregation strategies to represent a spectrum of complexity and theoretical guarantees. The random baseline serves as a control condition, enabling assessment of whether more sophisticated methods provide meaningful improvements. The Borda count was chosen for its simplicity and low computational overhead, and because it has been successfully applied to TCP by Mondal and Nasre (2021) in the Hansie framework. The Schulze method (Schulze, 2011), a Condorcet-consistent approach, was included to evaluate whether stronger theoretical properties (such as the majority winner criterion) translate to better TCP performance, despite its higher computational cost ($O(n^{3})$). This selection allows us to investigate the trade-off between computational efficiency and aggregation quality in the context of test case prioritization.
@@
For example, consider a RandomMixedOrder mixer combining ExeTimeOrder with a weight of 0.75 and RecentnessOrder with a weight of 0.25. When a test suite is provided, the mixer first sorts the test cases according to the rules of its sub-approaches. As a result, two permutations of the same test suite are produced: one ordered by prior execution times, and the other ordered by recentness. To produce the final order, the mixer repeatedly selects and runs the first unexecuted test case from one of the two lists at random, with the list based on execution-time having three times the probability of being chosen. This process continues until either all test cases in the suite have been executed or a failing case is encountered.